OpenStudy (anonymous):

Bobby bought 3 and 5/12 ft of red ribbon and 5 and 1/12 ft of yellow ribbon. Ribbon costs $1.20 a foot. How much did Bobby spend on ribbon? A. $9.60 B. $10.20 C. $10.32 D. $18.72

OpenStudy (anonymous):

I will give a medal!

OpenStudy (igreen):

For yellow.. 5 * 1.20 = $6.00 0.5 * 1.20 = $0.60 $6.00 + $0.60 = $6.60 So he spent $6.60 on yellow ribbons. For red.. 3 * 1.2 = $3.60 1.20 * 5/12 = $0.50 $3.60 + $0.50 = $4.10 So he spent $4.10 on red ribbons. $6.60 + $4.10 = $10.70 So he spent a total of $10.70 on ribbons.

OpenStudy (anonymous):

Thank you!